# TinkerXR: In-Situ, Reality-Aware CAD and 3D Printing Interface for Novices

### v3

@@ 3. DESIGN RATIONALE
When developing TinkerXR, several key design considerations were prioritized to ensure the system was both effective and intuitive for novice users. The main considerations included are user-friendly design and interaction, workspace awareness, and seamless integration:
3.1. User-Friendly Design and Interaction
To enhance the user experience in TinkerXR, we prioritize both simplicity in design and intuitive interaction. Our approach combines straightforward design logic with immersive user interactions to create an accessible and engaging environment for all users. The system’s design and interaction features are crafted to ensure that even novice users can seamlessly create and manipulate models, while advanced users benefit from an efficient design process.
@@ 3.1. User-Friendly Design and Interaction
Intuitive Interaction:
Interactivity in TinkerXR is driven by hand gestures and movement, recognized by Quest’s tracking system. These gestures allow users to manipulate objects directly, making the design process intuitive and engaging. Building upon the methods outlined in prior works (Kim et al., 2005; Weichel et al., 2014; Stemasov et al., 2024), TinkerXR lets users interact with the elements directly with their hands and without the need for external tools other than the headset. For instance, a pinching motion allows users to grab and move objects, while wrist rotation toggles features such as snap-to-grid, multi-selection, and uniform scaling. Moving the weak hand along a single global axis locks object manipulation to that axis, ensuring precision and control. Wrist-mounted menus provide an intuitive interface with buttons for quick navigation, including shortcuts for pinch gestures using fingers other than the index (e.g., select/deselect all). Visual and auditory feedback confirms actions, further enhancing the system’s usability and creating an immersive experience.
@@ 3.2. Workspace Awareness
Figure 2. Real-life workspace awareness. (a) Three cases for selecting the workspace in a scanned room. (b) Grid appearing on the selected workspace plane for the three corresponding cases. (c) Selecting the workspace based on a reference of a physical object. (d) Physical and digital objects rendered together with depth occlusion off (upper) versus depth occlusion on (lower).
[…]
Four photos labeled (a), (b), and two in (c). The first two photos show the first-person view of the user looking at a room with a table inside. The other two photos show three small decorative cacti on a table with a digital pyramid model behind them. In the image (a), three yellow rays come out from the user’s right hand, first hitting the floor where a text appearing above says ”select: floor”, second hitting the top of the table where a text appearing above says ”select: table”, and the third hitting the wall where a text appearing above says ”select: wall”. There is a white outline around the table. In image (b) the room is seen from the same angle as an image (a), but this time there are grids on top of the floor, table, and wall. In the first image of (c), the digital pyramid model partially blocks the view of the cacti even though it is placed behind them while in the second image of (c) the cacti are fully visible.
@@
It is also essential for users to understand their surroundings for context awareness while designing. While AR functionalities significantly aid in this area, models rendered in front of every world object can detract from the realistic nature we aim to achieve in the design task. By understanding the depth of objects in the environment, TinkerXR can occlude parts of a model that are behind physical objects, enhancing the realism and immersion of the design experience while elevating the user’s perception of the design as part of their real environment.
3.3. Seamless Integration
@@
In TinkerXR, we have consolidated design and fabrication functionalities into a single system, streamlining the entire workflow. Users can design their models and select printing options (e.g., layer height, infill pattern, support structure) directly within the AR environment, eliminating the need to transfer designs between separate tools. This integration simplifies the process by allowing users to handle all aspects of design and preparation within a unified interface, thereby reducing the complexity traditionally associated with using multiple software packages.
@@ 4. Implementation
Figure 3. User workflow. (a) Users begin by selecting a workspace from one of the flat surfaces in their room. (b, c) On the selected surface, users can create and manipulate primitive geometries using their hands while utilizing the tools provided by TinkerXR. In this example, the user designs a finger skateboard obstacle. (d) The final model is then placed in a virtual 3D printer twin where print slicing options can be adjusted before (e) fabricating on the remote physical printer. (f) Examples from the user study (Section 6) that demonstrate the design anywhere functionality, i.e., the flexibility to adapt to different environmental contexts.
[…]
@@ 4.2.1. Object Creation
Figure 4. Object creation and selection. (a, b, c) Poking gesture on buttons for creating primitive objects. (d) Highlighting by pointing hand towards an object and selecting it by pinch move in single selection mode. (e) Selecting objects in multiple selection mode.
[…]
Five photos labeled (a), (b), (c), (d), (e) show the first-person view of a user in front of a wall interacting with their hand and the digital objects. In images (a), (b), and (c) user have extended their weak hand to their eye level and looking at the inside of their wrist. There are 12 white square buttons on the wrist in a three-by-four grid. Seven of the buttons show primitive geometric objects, namely cube, cylinder, sphere, capsule, triangular prism, pyramid, and cone. In the image (a), the user is hovering his dominant hand index finger on their wrist. In the image (b), the user pokes the cube button with their dominant hand index finger and the button turns green and a cube appears above the weak hand. In the image (c), the user does the same with a sphere next to the cube. In image (d), user does a pinch motion while their dominant hand hand is directed toward the created cube and their weak hand is oriented palm-down. The cube’s color is changed from blue to cyan and a white grid is appearing on the cube. Similarly, in the image (e), the user does the same pinch motion for the sphere, this time with the weak hand oriented palm-up. Both of the objects turn to cyan, with white grid appearing on the sphere instead of the cube.
@@ 4.2.4. Object Manipulation
Figure 5. Moving, rotating, resizing objects. (a) Moving objects around the world space with the movement handle. (b) Rotating objects by twisting the rotation handle around its corresponding axis. (c) Resizing objects with the corner handles.
[…]
Six photos labeled (a), (b), (c) each label corresponding to two photos show a cube being manipulated with their dominant hand from the first-person view of the user in front of a wall. The first pair of pictures labeled (a) shows the user pinching the movement handle on top of the cube in the first picture and dragging their dominant hand and the object to the right in the second picture. The second pair of pictures labeled (b) shows the user pinching the rotation handle next to the cube in the first picture and rotating their dominant hand wrist and the cube in the second picture. The third pair of pictures labeled (c) shows the user pinching the corner handle at the top right corner of the cube in the first picture and moving their hand to the right and the cube stretching to the right in the second picture.
@@
A movement handle, i.e., gizmo, above the objects follows the user’s pinch gesture, which allows for freehand movement (Figure 5a).
@@
Three axis-aligned rotation handles allow rotation around the x, y, and z axes. Grasping a handle with a pinch gesture rotates objects like a knob (Figure 5b).
@@
Corner handles at each vertex of the Object Manipulation Box enable resizing. Adjusting a handle moves it freely while the opposite handle remains fixed, which ensures precise control (Figure 5c).
Uniform scaling activates when the weak hand is sideways (Figure 6a, 6b). A guiding line extends from the box center through the selected corner, ensuring accurate scaling along this line. Multiple selections allow synchronized scaling, while free scaling does not preserve aspect ratios (Figure 6c).
Figure 6. Uniform vs. non-uniform scaling. (a, b) Resizing objects with uniform scaling mode. (c) In free scaling mode.
[…]
Three photos labeled (a), (b), and (c) show a first-person view of a user looking at a wall and interacting with a complex 3D abstract design. In images (a) and (b), the user has their weak hand’s palm looking right and they apply a pinching motion to the corner handle on the upper right side of the manipulation box. A pink ray is seen coming from the middle and going through the selected corner handle. The user brings their right hand in the second picture closer to the center of the manipulation box, resulting in a smaller object with the same aspect ratio. Image (c) shows the user dragging the same corner handle with their weak hand palm looking down. The resulting object is compressed vertically.
@@
When any of the elements for manipulating objects are held with the dominant hand with a pinch gesture, the orientation of the weak hand dictates the toggle between free movement and snap-to-grid modes instead of single and multiple selection modes. Free movement is enabled when the weak hand’s palm faces upward, while snap-to-grid is activated when the palm faces downward. The functionality adapts to movement, rotation, and scaling operations.
Figure 7. Manipulating objects with snap-to-grid mode for (a) movement, (b) rotation, and (c) resizing.
[…]
Three photos labeled (a), (b), and (c) show the first-person view of the user looking at a wall with a vertical grid on the wall, holding their weak hand palm looking up, and manipulating a cube object. In image (a) user does a pinch motion to the movement handle on top of the cube. There are arrows appearing perpendicular to the sides of the cube with uniform steps. Similarly, in image (c), the same arrows are seen but for the scaling handle on the corner of the cube. In image (b), the user applies a pinch motion to the rotation handle on top of the cube and a circle with 15 degrees of steps parallel to the top side of the cube is seen.
For movement, when the snap-to-grid feature is activated, the Object Manipulation Box will align with the grid lines on the selected workspace (Figure 7a). In some cases, both parallel sides of an object may not be able to snap to the grid lines due to the distance between them. In such situations, the system determines which side should snap to a grid line based on the following logic:
@@
For rotation, when the snapping option is activated, the rotation handle adheres to a 15-degree increment from its initial position (Figure 7b). This 15-degree step is chosen to provide a balance between precision and ease of use, ensuring that users can achieve accurate rotations without making the process cumbersome.
For scaling, when the snap-to-grid feature is activated, the corner handle aligns with the nearest grid point within the scene space to the pinch location (Figure 7c). The objects are scaled in the same logic as the free movement mode, where the opposite corner acts as an anchor.
@@ 4.2.5. Axis Locking
The axis lock tool in TinkerXR allows users to easily constrain object manipulation to specific axes or planes, enhancing precision during design. Users can activate the axis lock tool by performing an index finger pinch gesture with their weak hand. Moving their hand in the desired global direction triggers a visual cue in the form of a colored edge (Figure 8a) or a side (if moved diagonally) of the axis lock tool, representing the active axis or plane of manipulation. This interaction ensures that users can limit movement or resizing action along a single axis or within a plane (Figure 8b and 8c) without needing to interact with additional buttons or menus. A single tap using the pinch gesture with their weak hand resets the axis lock, allowing users to resume unconstrained manipulation.
Figure 8. Axis locking. Users can (a) initiate the axis lock by pinching with their weak hand and moving it in the desired direction. Useful for limiting the (b) movement and (c) resizing of the objects.
[…]
@@ 4.2.7. Combining Objects
Users can combine and turn selected objects into one uniform object with a button on the weak hand wrist. The combined object will turn into the color of the last solid object chosen. The combining operation works with solid and hole objects, removing the hole objects after performing a cut on the solid objects through the intersection between the solid and hole objects (Figure 9b and 9c).
Figure 9. Multi-object operations. Users can create (a) solid and hole objects. (b, c) The combining operation creates a cut in the intersection between the solid and hole object.
[…]
Three photos labeled (a), (b), and (c) show a first-person view of a user looking at a wall. In the image (a) a solid cube and a hole sphere intersecting with the top corner of the cube is seen. The hole (or negative) sphere is seen with red color and with dotted transparent skin. In image (b) the weak hand wrist is visible while the cube and the sphere are selected and colored cyan. The user is about to poke a button named ”combine” with their dominant-hand index finger. In the image (c), the combination of the objects is visible where there is a round cut on the upper corner of the cube where the sphere used to intersect the cube.
@@ 4.2.8. Reference Objects
Figure 10. Reference objects. Users can (a) create reference objects via the dominant-hand menu and (b) use them to represent physical items. (c) On-object grids are visualized for precise design. (d) The example shows a plant pot hanger designed around a reference pot.
[…]
4.2.9. Ruler Tool
Users can create a virtual ruler by pressing a button on their weak hand wrist. This simple ruler, equipped with two interactable edge handles (similar to scaling handles), provides real-world scale measurements. This ruler can be used for precise alignment and measuring distances between objects within the real and virtual space (Figure 11a).
Figure 11. In-context measurement. Users can (a) measure the length of the objects around them with the ruler tool. For parametric resizing, users (b) select the handles on the edges of the object manipulation box, which (c) activates a numerical keyboard on the weak hand for (d) inputting the desired length alongside the edges.
[…]
Four images labeled (a), (b), (c), and (d) show a user’s first-person view using the system. In picture (a), user is using their weak and dominant hand to pinch both ends of a digital line where in the middle of the line it says ”20.00 cm”. The user is measuring out the length of a game console (Nintendo Switch). In pictures (b), (c), and (d) user is manipulating a digital cube model. In the picture (b) user is applying a pinch motion to a small black handle on the middle of the edge of the cube. In picture (c), a numerical keyboard interface appears on the weak hand in direction of the palm and it says ”10 cm” in the input field of the keyboard interface. In picture (d), the user has inputted ”20 cm” as the input and clicks the confirmation button. The cube’s width is doubled.
4.2.10. Parametric Resizing
In addition to free, snap-to-grid, and uniform scaling operations, users can also adjust object sizes through precise numerical length inputs. To use this feature, users first select the edge of the Object Manipulation Box they wish to resize (Figure 11b) using a pinch gesture. Then, a numerical keyboard appears on the user’s weak hand (Figure 11c), allowing them to enter the desired real-world length for the selected edge. The object scales proportionally along the direction of the chosen edge based on this input (Figure 11d). This method can be effectively combined with the ruler tool to ensure accurate design according to specific measurements.
@@ 4.3. Fabrication Features
After the target model has been created, the user can export it to an STL file. This will allow users to access it from other CAD tools and enable them to print their designs. This feature not only accelerates and simplifies the fabrication process but also enables users to quickly fabricate a draft of their design to assess whether it will be printed with the fidelity they anticipate.
@@
Figure 12. Transition to fabrication. Virtual printer twin is used as a placeholder for exporting the model for a target printer or conducting a quick size check. Users can (a, b) drag and drop the models they create in the printer, (c) modify the print orientation, and (d) instantiate physical printing.
[…]
Four images labeled (a), (b), (c), and (d) show a user’s first-person view looking on top of a table where there is a horizontal grid appearing on the table and a virtual 3D printer twin and a model of a wall hook is visible. The 3D printer twin is seen as a transparent rectangular prism with white edges and a grid on the bottom plate. The name of the printer (Creality Ender-3 V3 KE) is written on the bottom of the printer. In the picture (a), the user is moving the model with the movement handle grabbed with a pinch gesture with the dominant hand, dragging it to the printer twin. A white line grid appears on the printer twin. In picture (b), the model is placed in the middle of the printer twin and it says ”model printed successfully” in green color on top of the printer twin. On the right side of the printer twin, two input fields and two buttons are visible. The first input field is labeled ”enter sever IP…” and the second input field is labeled ”enter printer IP…”. The first button is labeled ”Start print” and the second button is labeled ”Slicer settings”. In the picture (c), user is rotating the model with a pinch gesture applied to the rotation handle on top of the model. In picture (d), a progress bar is seen on top of the printer twin and a text is written as ”0%” in the middle. The input fields contain the partially blurred IP addresses of the server and the printer and the buttons are labeled ”pause print” and ”abort print”.
Figure 13. Preparing for printing and slicing. (a) Users can select the slicing options for their print. They initiate slicing (b) and (c) view the sliced preview of the model, with outer walls in gray and support structures in cyan. (d) If a model exceeds the printer’s build area, a ”too large for printer” indicator appears above the printer twin.
[…]
The low-poly Stanford bunny is placed inside the virtual printer twin with slicing options displayed next to it.
4.3.1. Spatial Virtual Twin of Printer
A printer twin is a unique object that cannot be selected in the multiple selection mode or cannot be merged with other objects. This virtual printer object helps the user estimate how the object will be printed inside the real-world 3D printer. The size of the virtual printer twin corresponds to the real-world printer’s build area. This virtual printer’s size can be selected from the predefined printers or set manually. To start the printing process, the user has to drag and drop the current model into the printer twin. If neither the server IP nor printer IP is given on the printer twin, it will save the current model locally into an STL file. If the slicer server is connected to the headset, the user can send their STL files to the server to create a G-code file for 3D printers. If the printer is also connected to the server, then the printing process will begin automatically.

### v4

@@ 3. DESIGN RATIONALE
When developing TinkerXR, several key design considerations were prioritized to ensure the system was effective for novice users. The main considerations included are novice-friendly design and interaction, workspace awareness, and seamless integration:
3.1. Novice-Friendly Design and Interaction
To enhance the user experience in TinkerXR, we prioritize both simplicity in design and interaction. Our approach combines straightforward design logic with immersive user interactions to create an accessible and engaging environment for all users. The system’s design and interaction features are crafted to ensure that even novice users can seamlessly create and manipulate models, while advanced users benefit from an efficient design process.
@@ 3.1. Novice-Friendly Design and Interaction
Natural Interaction:
Interactivity in TinkerXR is driven by hand gestures and movement, recognized by Quest’s tracking system. These gestures allow users to manipulate objects directly. Building upon the methods outlined in prior works (Kim et al., 2005; Weichel et al., 2014; Stemasov et al., 2024), TinkerXR lets users interact with the elements directly with their hands and without the need for external tools other than the headset. For instance, a pinching motion allows users to grab and move objects, while wrist rotation of the weak hand toggles features such as snap-to-grid, multi-selection, and uniform scaling. Moving the weak hand along global axes (x, y, z) toggles axis and plane locks for much needed precision while manipulating objects. The wrist-mounted menus provide an interface with buttons for quick navigation of available operations and settings. To increase design productivity, there are also shortcuts created that utilize pinch gestures with middle and ring finger for selection operations that do not require pointing at an object, which are the selection and deselecting of all objects created. Visual and auditory feedback confirms actions to further enhance the system’s usability and create an immersive experience.
@@ 3.2. Workspace Awareness
Figure 3. User workflow. (a) Users begin by selecting a workspace from one of the flat surfaces in their room. (b, c) On the selected surface, users can create and manipulate primitive geometries using their hands while utilizing the tools provided by TinkerXR. In this example, the user designs a finger skateboard obstacle. (d) The final model is then placed in a virtual 3D printer twin where print slicing options can be adjusted before (e) fabricating on the remote physical printer. (f) Examples from the user study (Section 6) that demonstrate the design anywhere functionality, i.e., the flexibility to adapt to different environmental contexts. :
@@
It is also essential for users to understand their surroundings for context awareness while designing. While AR functionalities significantly aid in this area, models rendered in front of every world object can detract from the realistic nature we aim to achieve in the design task. By understanding the depth of objects in the environment, TinkerXR can occlude parts of a model that are behind physical objects. This enhances the realism and immersion of the design experience, while elevating user perception of the design as part of their real environment.
3.3. Unified Design-to-Fabrication Workflow
@@
In TinkerXR, we have consolidated design and fabrication functionalities into a single system, streamlining the entire workflow. Users can design their models and select print settings (e.g., layer height, infill pattern, support structure) directly within the AR environment, eliminating the need to transfer designs between separate tools. This simplifies the process by allowing users to handle all aspects of design and preparation within a unified interface, thus reducing the complexity traditionally associated with the use of multiple software packages.
@@
Figure 4. Object creation and selection. (a, b, c) Poking gesture on buttons for creating primitive objects. (d) Highlighting by pointing hand towards an object and selecting it by pinch move in single selection mode. (e) Selecting objects in multiple selection mode.:
@@ 4.2.4. Object Manipulation
Figure 5. Different ways of manipulating objects. Users can perform the following main actions: Moving objects around the world space with the movement handle; rotating objects by twisting the rotation handle around its corresponding axis; and resizing objects by dragging the corner handles. The orientation of the weak-hand toggles manipulation modes between free movement (palm-down), uniform scaling (palm-sideways), and snapping-to-grid (palm-up).:
@@
A movement handle, i.e., gizmo, above the objects follows the user’s pinch gesture, which allows for freehand movement (Figure 5).
@@
Three axis-aligned rotation handles allow rotation around the x, y, and z axes. Grasping a handle with a pinch gesture rotates objects like a knob (Figure 5).
@@
Corner handles at each vertex of the Object Manipulation Box enable resizing. Adjusting a handle moves it freely while the opposite handle remains fixed, which ensures precise control (Figure 5).
Uniform scaling
Gets activated when the weak hand is sideways (Figure 5). A guiding line extends from the box center through the selected corner, ensuring accurate scaling along this line. Multiple selections allow synchronized scaling, while free scaling does not preserve aspect ratios.
@@
When any of the elements for manipulating objects are held with the dominant hand with a pinch gesture, the orientation of the weak hand dictates the toggle between free movement and snap-to-grid modes instead of single and multiple selection modes. Free movement is enabled when the weak hand’s palm faces upward, while snap-to-grid is activated when the palm faces downward. The functionality adapts to movement, rotation, and scaling operations (Figure 5).
For movement, when the snap-to-grid feature is activated, the Object Manipulation Box will align with the grid lines on the selected workspace. In some cases, both parallel sides of an object may not be able to snap to the grid lines due to the distance between them. In such situations, the system determines which side should snap to a grid line based on the following logic:
@@
Figure 6. Axis locking. Users can (a) initiate the axis lock by pinching with their weak hand and moving it in the desired direction. This helps (b) restrict the movement and (c) resize the objects.:
For rotation, when the snapping option is enabled, the rotation handle aligns with 15-degree increments from its initial position by default. Smaller increment values allow for greater rotational flexibility, though they may reduce precision in achieving the desired orientation due to increased likelihood of overshooting or undershooting, especially at finer resolutions such as 1-degree steps.
For scaling, when the snap-to-grid feature is activated, the corner handle aligns with the nearest grid point within the scene space to the pinch location. The objects are scaled in the same logic as in the free-motion mode, where the opposite corner acts as an anchor. We note that the cell size of the grid for movement and scaling, as well as the snapping degrees for rotation are adjustable through settings.
Figure 7. Reference objects. Users can (a) create reference objects via a button and (b) use them to represent physical items. (c) On-object grids can be created for alignment on the reference object for an object-focused design. (d) The example shows a plant pot hanger designed around a reference pot.:
@@ 4.2.5. Axis Locking
The axis lock tool in TinkerXR allows users to easily constrain object manipulation to specific axes or planes. While this is a fundamental operation, our approach to selection logic sets this mechanism apart from others. Users can activate the axis lock tool by performing an index finger pinch gesture with their weak hand. Moving their hand in the desired global direction triggers a visual cue in the form of a colored edge (Figure 6a) or a side (if moved diagonally) of the axis lock tool, representing the active axis or plane for manipulation. This interaction ensures that users can limit movement or resizing action along a single axis or within a plane (Figure 6b and 6c) without needing to interact with additional buttons or menus.
A single tap using the pinch gesture with the weak hand resets the axis lock and can start moving in all directions again. We chose this to allow users to resume unconstrained manipulation quickly. In the development phase, we tested axis-specific handles around the Object Manipulation Box, but it was apparent that these handles drastically crowded the area around the selected objects, which made the objects harder to see and interact with.
@@ 4.2.7. Combining Objects
Users can combine and turn selected objects into one uniform object with a button on the weak hand wrist. The combined object will turn into the color of the last solid object chosen. The combining operation works with solid and hole objects, removing the hole objects after performing a cut on the solid objects through the intersection between the solid and hole objects (Figure 8b-d).
Figure 8. Multi-object operations and in-context measurement. Users can create (a) solid and hole objects. (b, c, d) The combining operation creates a cut in the intersection between the solid and hole object. Users can (e) measure the length of the objects around them with the ruler tool. For parametric resizing, users (f) select the handles on the edges of the object manipulation box, which (g) activates a numerical keyboard on the weak hand for (h) inputting the desired length alongside the edges.:
4.2.8. Ruler Tool
Users can create a virtual ruler by pressing a button on their weak hand wrist. This simple ruler, equipped with two interactable edge handles (similar to scaling handles), provides real-world scale measurements. This ruler can be used for precise alignment and measuring distances between objects within the real and virtual space (Figure 8e).
4.2.9. Parametric Resizing
In addition to resizing with free movement, snapping-to-grid, and uniform scaling operations, users can also adjust object sizes through precise numerical length inputs. To use this feature, users first select the edge of the Object Manipulation Box they wish to resize (Figure 8f) using a pinch gesture. Then, a numerical keyboard appears on the user’s weak hand (Figure 8g), allowing them to enter the desired real-world length for the selected edge. The object scales proportionally along the direction of the chosen edge based on this input (Figure 8h). This method can be effectively combined with the ruler tool to ensure accurate design according to specific measurements.
Figure 9. Transition to fabrication. Virtual 3D printer twin is used as a placeholder for exporting the model for a target printer or conducting a quick size check. Users can (a, b) drag and drop the models they create in the printer, (c) modify the print orientation, and (d) instantiate physical printing.:
@@ 4.3. Fabrication Features
After the target model has been created, the user can export it to an STL file. This will allow users to access it from other CAD tools and enable them to 3D print their designs. This feature not only accelerates and simplifies the fabrication process but also enables users to quickly fabricate a draft of their design to assess whether it will be printed with the fidelity they anticipate.
@@
Figure 10. Preparing for 3D printing and slicing. (a) Users can select the slicing options for their print. They initiate slicing (b) and (c) view the sliced preview of the model, with outer walls in gray and support structures in cyan. (d) If a model exceeds the 3D printer’s build area, a ”too large for printer” indicator appears above the printer twin.:
4.3.1. Spatial Virtual Twin of 3D Printer
A printer twin is a unique object that cannot be selected in the multiple selection mode or cannot be merged with other objects. This virtual printer object helps the user estimate how the object will be printed inside the real-world 3D printer. The size of the virtual printer twin corresponds to the real-world 3D printer’s build area. This virtual printer’s size can be selected from the predefined printers or set manually. To start the 3D printing process, the user has to drag and drop the current model into the printer twin. If neither the server IP nor the printer IP is given on the printer twin, it will save the current model locally into an STL file. If the slicer server is connected to the headset, the user can send their STL files to the server to create a G-code file for 3D printers. If the printer is also connected to the server, then the 3D printing process will begin automatically.